$SPAD/src/input erf.input

Waldek Hebisch December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)spool erf.output
)set message test on
)set message auto off
)clear all
--S 1 of 4
i:=integrate(exp(-a^2*p^2),p=b..c)
--R
--R                              +---+
--R        (erf(a c) - erf(a b))\|%pi
--R   (1)  ---------------------------
--R                     2a
--R                  Type: Union(f1: OrderedCompletion(Expression(Integer)),...)
--E 1
--S 2 of 4
eval(i,[erf(a*c)=1,erf(a*b)=0])
--R
--R         +---+
--R        \|%pi
--R   (2)  ------
--R          2a
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 4
intrule:=rule integral(exp(-a*p^2),p) == sqrt(%pi)/sqrt(a)
--R
--R           p       2          +---+
--R         ++    - %I a        \|%pi
--R   (3)   |   %e      d%I  == ------
--R        ++                     +-+
--R                              \|a
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 3
--S 4 of 4
intrule integral(exp(-a*p^2),p)
--R
--R         +---+
--R        \|%pi
--R   (4)  ------
--R          +-+
--R         \|a
--R                                                    Type: Expression(Integer)
--E 4

)spool )lisp (bye) ———-

4

References^{[1] nothing}

5